Solid sodium metal is added to liquid water to produce hydrogen gas and aqueous sodium hydroxide.

1. Write the balanced reaction.

2. If 24.5 g of water were reacted, how many moles is that?.

3. If 8.95 moles of hydrogen gas were produced, how many germs is that?.

4 If 3.45 moles of sodium hydroxide are produced, how many moles of sodium metal reacted?

5. What type of reaction is this?.


1
Expert's answer
2021-03-03T01:57:55-0500

1.)2Na+ 2H2O= 2NaOH + H2

2.) 24.5g of water is equivalent to

=24.5/18=1.36moles

3.)mass="Moles\u00d7 rfm"

= 8.95× 2= 17.9g

4) mole ratio 1:1

There 3.45mole of Na reacted

5)redox reaction